OpenStudy (anonymous):

(MEDAL & FAN) Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3

OpenStudy (anonymous):

it is x-y=7

OpenStudy (anonymous):

A. x = (7 – x) + 3 B. x – (y + 3) = 7 C. x – y = y + 3 D. (y + 3) – y = 7 is my choices

OpenStudy (anonymous):

everything blued is a -

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so do i use a b c or d i am lost

OpenStudy (michele_laino):

please, you have to use D

OpenStudy (anonymous):

i will try it thank you

OpenStudy (anonymous):

i will tell you right or wrong

OpenStudy (michele_laino):

yes, I'm waiting

OpenStudy (anonymous):

can you help me with 2 more that way i can check

OpenStudy (michele_laino):

Yes!

OpenStudy (anonymous):

What is the solution to the system of equations? i need 2 number for slope intercept form http://static.k12.com/bank_packages/files/media/mathml_c54e6357d10d28554e1afa49011c16f6dda9011e_1.gif

OpenStudy (michele_laino):

please, if you apply the substitution method for solving your system, you find, the solution below: x=-3, and y=-2 from the second equation of the system I get the value of x, as: \[x=\frac{ 4+5y }{ 2 }\]substituting that value of x in the first equation of your system, you get: \[5*\frac{ 4+5y }{ 2 }-3y=-9\] please check my statements

OpenStudy (anonymous):

ok thank you last one now

OpenStudy (michele_laino):

Ok! go please!

OpenStudy (anonymous):

https://storage.googleapis.com/clipular-clips/4530423126294528/6601633364443136?GoogleAccessId=clipular2012 @appspot.gserviceaccount.com&Expires=1416687227&Signature=OETLMAzJxS3OCL9yby9A3bGBkiWezauhvQ8Aesm8w7NqNO1MXCDKp5W7jWoXSTuXfThzHL9Djka6SwLOw%2B3au3fOe02b1Wm%2BwpfuJNHUVn85mxa2LL5muOAV%2BKydb5mbGxeir2C51G5CK7zs3HTogeA1iFR4NezBjh314u2UCho%3D&md5=6d756b6faec5425b21457c2a14d39465&NOT_A_LINK

OpenStudy (michele_laino):

Sorry I can't see the page

OpenStudy (anonymous):

highlight the whole thing not click the blue one

OpenStudy (anonymous):

then copy and paste at top of serch engine

OpenStudy (anonymous):

nvm lol now it works

OpenStudy (anonymous):

http://static.k12.com/bank_packages/files/media/mathml_465c125993163872ea9b275109dbbbbd92a018c3_1.gif The system of equations is coincident. What are the missing values?

OpenStudy (anonymous):

https://storage.googleapis.com/clipular-clips/4530423126294528/6601633364443136?GoogleAccessId=clipular2012 @appspot.gserviceaccount.com&Expires=1416687435&Signature=QwBWJm653xqFLlvsGLkp21THgCBn3OfHGEQqijw%2BdEpC8Dn914V5H8d1hnzqCCf3i3fh3nZNMgAB6mV3GJet2d0EDa7DpsSc%2B9qSqNcwXSo5NJTkHiWLJ2RCLr9%2FnNkigXAUz%2BgPwwZjvkgmsg0gVT0jgS%2FT6FdZef0Vrp%2FpUGA%3D&md5=6d756b6faec5425b21457c2a14d39465&NOT_A_LINK

OpenStudy (anonymous):

lol it just sucks

OpenStudy (michele_laino):

why you say that the system is coincident? I don't understand

OpenStudy (michele_laino):

do you know what is the meanig of system coincident?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

Please tell me!

OpenStudy (michele_laino):

@boygame20

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://www.mathnstuff.com/math/spoken/here/1words/c/c22.htm here this have all the meaning

OpenStudy (michele_laino):

Ok! I understand!

OpenStudy (michele_laino):

I think that the second equation is: 9x+3y=21, so the missing values,are: 3 and 21.

OpenStudy (michele_laino):

the coefficients of the second equation are the same of the first equation, namely 3x+y=7, multiplied by 3

OpenStudy (anonymous):

ok well i try them

OpenStudy (anonymous):

OMG I JUST DONE IT the ones i did made me fail will you help me agian

OpenStudy (anonymous):

yours were all right

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x – 9y = 0 –x + 3y = –3 A. 0 B. 2 C. infinitely many D. 1

OpenStudy (michele_laino):

Ok! I help you, if you have other exercise, please tell me!

OpenStudy (anonymous):

no i am good this mandatory i have to pass with 80%

OpenStudy (michele_laino):

I think A. as you can check using the substitution method

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you certain i not good with that method that why i failed

OpenStudy (michele_laino):

if you multiply both sides of the second equation, you get: 3x-9y=-3, here the left side is equals to the left side of the first equation, nevertheless the right sides of the two equations are not equals, so it is impossible. That's because we have no solutions

OpenStudy (anonymous):

ok it is the one you did before kinda

OpenStudy (anonymous):

What is the solution to the system of equations? http://static.k12.com/bank_packages/files/media/mathml_d676641518b9981930709d538c634ae09a39918d_1.gif

OpenStudy (anonymous):

ok but next one please thank you though i get it now

OpenStudy (haleyelizabeth2017):

Are you needing to show it being substituted?

OpenStudy (anonymous):

i need it for my school i need it in slope intercept form

OpenStudy (haleyelizabeth2017):

oh! Okay :D slope intercept form is simple......y=mx+b, correct?

OpenStudy (anonymous):

yes

OpenStudy (haleyelizabeth2017):

Are we working on the two equations? Or is that for something different?

OpenStudy (anonymous):

i need the x and y that makes those true

OpenStudy (haleyelizabeth2017):

oh! so we will just use substitution to find that.....okay?

OpenStudy (haleyelizabeth2017):

Which equation do we need to turn into slope-intercept form?

OpenStudy (anonymous):

yes i just need the anser in that form (sorry for my spelling)

OpenStudy (anonymous):

none i just need the x and y i got confused

OpenStudy (michele_laino):

from the second equation I get: \[x=-4+2y\] inserting that value for x into the first equation, we have: \[2(-4+2y)+y=-3\] from which, y=1. substituting y=1, into the expression for x, above, we have: x=-2, please check my statements

OpenStudy (anonymous):

i got another 7 qustions after this

OpenStudy (haleyelizabeth2017):

oh.....okay

OpenStudy (haleyelizabeth2017):

I will let @Michele_Laino help you out because I have math of my own to do today as well :( It's difficult so I need to focus...sorry!

OpenStudy (michele_laino):

what is your new Mathematical question, please?

OpenStudy (anonymous):

still this one he never got me past it http://static.k12.com/bank_packages/files/media/mathml_d676641518b9981930709d538c634ae09a39918d_1.gif

OpenStudy (michele_laino):

from the second equation, I get: \[x=-4+2y\] now I'going to insert that expression for x, into the first equation, so I can write: \[2*(-4+2y)+y=-3\] please simplify this last equation, and find y

OpenStudy (anonymous):

can you try i am trying to work number 8 your working 3

OpenStudy (michele_laino):

here are the steps: \[-8+4y+y=-3\] \[-8+5y=-3\] \[5y=-3+8\] \[5y=5\] \[y=1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

umm do you know x please i in a rush i only got 30 mins left

OpenStudy (michele_laino):

now insert y=1, into the expression for x, namely \[x=-4+2y\] and write your result for x, please

OpenStudy (anonymous):

so x=-2?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ok next

OpenStudy (anonymous):

In a recent golf match, Desmond’s score was 3 less than George’s score. Their combined scores totaled 187. Let d represent Desmond's score and g represent George's score. Which pair of equations can be used to determine their scores? A. g = d – 3 d + g = 187 B. d + g = 187 d – g = 3 C. d = g – 3 d + g = 187 D. d + g = 3 187 + d = g

OpenStudy (anonymous):

opss wrong one

OpenStudy (anonymous):

here let me get 6

OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. x + y = y + 5 B. x = (x – 6) + 5 C. x + (y + 5) = 6 D. (y + 5) + y = 6

OpenStudy (michele_laino):

for the golf match question, the answer is C.

OpenStudy (anonymous):

ok i got 2 more for you

OpenStudy (michele_laino):

the right answer is D.

OpenStudy (anonymous):

:/ i need to pass no joke please

OpenStudy (anonymous):

i put d

OpenStudy (michele_laino):

I'm a serious boy , don't worry!

OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values? http://static.k12.com/bank_packages/files/media/mathml_56c7fd0a18144c3c3e4379c77c9b6e184a7b51bb_1.gif http://static.k12.com/bank_packages/files/media/mathml_e8e38041922035857054d4ea27e296efc8c85734_1.gif and ok i need the black's filled in

OpenStudy (anonymous):

blank

OpenStudy (michele_laino):

In one page I see only 5x, in the second page I see the equation x+3y=8

OpenStudy (anonymous):

oh no worry i will get my mom to try it agian thank you if i fail i need you agian

OpenStudy (michele_laino):

Ok!

OpenStudy (michele_laino):

Sorry, what is the meaning of inconsistent system

OpenStudy (anonymous):

I PASSED THANK YOU